What is the nth term of 1, -2, 4, -8,16
Found 2 solutions by Sir226, Edwin McCravy:
Answer by Sir226(11)

1,-2,4,-8,16
t2/t1=-2/1
r=-2
Tn=ar^n-1
=1÷ -2^n-1

=-2^n-1

Answer by Edwin McCravy(20055)

 
1, -2, 4, -8,16

1st term = 1 = 20
2nd term = 2 = 21
3rd term = 4 = 22
4th term = 8 = 23
5th term = 16 = 24

Notice that the exponent of 2 is 1 less than
the number of term.

nth term = 2n-1 ignoring sign.

But then you need the sign-changer factor (-1)n-1.	

nth term = (-1)n-12n-1

an = (-1)n-12n-1


Edwin